A block of mass m1, placed on an inclined smooth plane is connected by a light string passing over a smooth pulley to mass m2, which moves vertically downwards. Find the tension in the string.


1
Expert's answer
2021-10-24T18:20:54-0400

Draw the free-body diagram:



Apply Newton's second law and assume the motion occurs at constant acceleration:


"1:-m_1a=-T+m_1g\cos\theta,\\
2: T-m_2g=-m_2a.\\\space\\

T=m_2(g-a),\\
-m_1a=-m_2g+m_2a+m_1g\cos\theta,\\
a(m_2+m_1)=m_2g-m_1g\cos\theta,\\\space\\
a=\frac{m_2g-m_1g\cos\theta}{m_2+m_1},\\\space\\
T=m_2\bigg(g-\frac{m_2g-m_1g\cos\theta}{m_2+m_1}\bigg)."